Calc help!
ta (16 x sinly dxdy) Il the 000 Coth - 8 Wa 2 nel (94+)? +(74)* +(-98–3t) th

coth (Pi)

its three sperate equations, stuck in the first one and the last one. the middle one is 1.6 because of the infinite N
Answer #1

\text{Let us evaluate the integral first}\\ \text{Suppose }I=16\int_0^\pi\int_0^1 x^2\sin y dxdy\\ \text{ Since the limits are constant, hence splitting them into two single integration}\\ \implies I=16\int_0^\pi \sin y dy\times \int_0^1x^2dx\\ \implies I=16\left ( -\cos y \right )_0^\pi\times \left ( \frac{x^3}{3} \right )_0^1\\ \implies I=16\times 2\times \frac{1}{3}=\color{red}\frac{32}{3}

Now evaluating the sum of a more general infinite series as follows:

In your case, you have a=1. Put a=1 in the last result;

\sum_{n=-\infty}^\infty \frac{1}{1+n^2}=\color{red}\pi\cot h\pi

Hence the first problem reduces to :

16\int_0^\pi\int_0^1x^2\sin y\ dxdy\left [\frac{\sum_{n=-\infty}^\infty \frac{1}{1+n^2}}{\cot h\pi}\right ]=\color{red}\frac{32}{3}\times\frac{\pi\cot h\pi}{\cot h\pi}=\frac{32\pi}{3}